OpenStudy (anonymous):

Determine the slope of the line with coordinates (-4,0) and (0,2)

OpenStudy (saifoo.khan):

\[slope = \frac{(y_2-y_1)}{(x_2-x_1)}\]

OpenStudy (saifoo.khan):

1/2

OpenStudy (anonymous):

right 4 up 2 slope is 2/4 = 1/2